# Non-Stationary Lipschitz Bandits

### v1

@@ 1 Introduction
Many practical applications involve decision-making over a continuous space of actions (or arms) where the environment evolves over time. For example, in dynamic pricing, the space of possible prices is inherently continuous, and market dynamics may shift abruptly. In such setting, it is both natural and crucial to leverage smoothness in the reward function to generalize across similar actions, while also remaining adaptive to temporal changes. However, most existing works on non-stationary bandits focus on discrete action spaces and overlook the structure present in continuous domains.
@@
The non-stationary bandit problem extends the classical stochastic multi-armed bandit setting (Slivkins et al., 2019; Lattimore and Szepesvári, 2020) by allowing the mean reward functions to change over time. Over the past two decades, several frameworks have been proposed to quantify non-stationarity. Common approaches quantify the number of changes $L_{T}$ of any arm’s mean reward over $T$ rounds (Auer et al., 2002b; Garivier and Moulines, 2011), or their total variation $V_{T}$ in the reward functions (Besbes et al., 2014; Cheung et al., 2019). Algorithms that are built upon these approaches achieve dynamic regret bounds of order $\sqrt{L_{T}T}$ or $V_{T}^{1/3}T^{2/3}$ respectively. Importantly, they all assume prior knowledge of the degree of non-stationarity $L_{T}$ or $V_{T}$. It remained an open question whether such rates could be achieved without knowing $L_{T}$ or $V_{T}$, until a line of work answered this affirmatively (Auer et al., 2019a, b; Chen et al., 2019; Wei and Luo, 2021).
@@
Recent works have introduced more refined measures of non-stationarity. In fact, the standard metrics $L_{T}$ and $V_{T}$ can be overly pessimistic: both of them may be arbitrarily large even when the best arm remains unchanged. To address this, Abbasi-Yadkori et al. (2023) designed an algorithm that tracks the number of times the best arm changes over $T$ rounds, $S_{T}$, and showed that it achieves $\sqrt{S_{T}T}$ for the dynamic regret, adaptively [^1] In all our work, ”adaptively” should be understood the same as ”without knowledge of the non-stationarity”.. The important work of Suk and Kpotufe (2022) proposed a finer notion of non-stationarity called significant shifts, which identifies only changes that meaningfully affect regret through large aggregate differences in reward. In the $K$-armed bandit setting, their algorithm achieves the optimal rate of $\sqrt{\tilde{L}_{T}T}$ where $\tilde{L}_{T}$ is the number of significant shifts over $T$ rounds, with $\tilde{L}_{T}\ll L_{T}$.
@@
In this work, we extend the ideas of Suk and Kpotufe (2022) to the setting of Lipschitz bandits, where the arm space is continuous (typically the interval $[0,1]$) and the reward functions satisfy a Lipschitz condition at each round. Despite the maturity of research in both non-stationary and Lipschitz bandits individually, their intersection remains unexplored. Our work is the first one to tackle the problem of non-stationary Lipschitz bandits.
@@ 1.1 Outline and contributions
Discretization level. In the stationary Lipschitz bandit setting, a common approach is to discretize the continuous arm space into a finite set of bins, thereby allowing the application of standard multi-armed bandit algorithms (Kleinberg, 2004). When the time horizon $T$ is known, the discretization level can be carefully calibrated, typically using $\mathcal{O}(T^{1/3})$ bins. This choice reflects the optimal balance between discretization error and the regret incurred by running a bandit algorithm over the discretized arms, leading to the minimax-optimal regret of $\mathcal{O}(T^{2/3})$. In the non-stationary setting, however, the lengths of locally stationary phases are unknown (and even not properly defined). Thus, direct horizon-dependent tuning becomes infeasible. This raises a key question: how to adaptively adjust the discretization level to match the unknown degree of non-stationarity in the environment?
Forced exploration via restarts and replays. In non-stationary bandits, detecting changes in the reward distribution requires frequent exploration of suboptimal arms. In the $K$-armed bandit setting, this is often handled through periodic restarts, where all $K$ arms are re-explored to detect shifts of the reward function. In the continuous setting, this strategy is significantly harder: not only must we decide when to restart, but also at what resolution to discretize the arm space during such restarts. Finer discretization enables more precise detection of changes but significantly increases the cost of exploration. Balancing this three-way trade-off is a central technical challenge tackled in our work.
This paper is organized as follows. In Section 2, we formally introduce the problem setting, and present our algorithm, MDBE, in Section 3. Our main results are stated in Section 4, followed by a high-level proof sketch of its regret analysis in Section 5. For completeness, the full pseudocode of MDBE is provided in Appendix B, and all detailed proofs can be found in Appendix D.
@@ 1.2 Related Work
Non-stationary bandits beyond $K$-armed. A growing body of work has extended non-stationary bandit models beyond the classical $K$-armed setting. These include linear and generalized linear bandits (Russac et al., 2019; Faury et al., 2021), convex bandits (Zhao et al., 2021; Wang, 2023), kernelized bandits (Hong et al., 2023; Iwazaki and Takeno, 2025; Cai and Scarlett, 2025), and more structured problems (Seznec et al., 2020; Vernade et al., 2020). Some works further explore temporal regularity assumptions, such as smooth or slowly varying reward functions (Slivkins and Upfal, 2008; Krishnamurthy and Gopalan, 2021; Jia et al., 2023). Across these settings, non-stationarity is typically quantified using either the number of global changes $L_{T}$ or the total variation $V_{T}$.
@@
Bandits with infinite arms. In Lipschitz bandits (Agrawal, 1995), the infinite arm space is tackled via discretization strategies, either fixed or adaptive, that exploit the Lipschitz continuity of the reward function (Kleinberg, 2004; Kleinberg et al., 2008; Bubeck et al., 2011b, a; Magureanu et al., 2014). More recently, adversarial formulations have been studied, where the goal is to minimize the regret in hindsight (Podimata and Slivkins, 2021; Kang et al., 2023). These differ fundamentally from our focus on dynamic regret in a stochastic setting. A separate line of work investigates infinite-armed bandits in the so-called reservoir model, where arm means are drawn independently from a fixed distribution (Wang et al., 2008; Carpentier and Valko, 2015; De Heide et al., 2021; Kim et al., 2022, 2024). Unlike Lipschitz bandits, these models assume no structural relationship between arms and therefore do not exploit similarity across them. Hence, direct comparisons to our setting do not apply.
@@
Tracking significant shifts. The idea of tracking significant shifts was first introduced in the $K$-armed setting by Suk and Kpotufe (2022), and has since been extended to contextual bandits (Suk and Kpotufe, 2023), preference-based settings (Suk and Agarwal, 2023; Buening and Saha, 2023; Suk and Agarwal, 2024), and smooth non-stationary models with temporal structure (Suk, 2024).
@@
The works closest to ours are Suk and Kpotufe (2023) and Suk and Kim (2025), both extending the framework of tracking significant shifts beyond the classical $K$-armed setting. The former considers contextual bandits where rewards are Lipschitz in the context space, but with a finite number of arms. After discretizing the context space, their algorithm effectively reduces to a finite-arm bandit problem at each context, allowing the use of techniques from the $K$-armed setting (Suk and Kpotufe, 2022) without requiring discretization of the arm space. The latter studies infinite-armed bandits under a reservoir model, where arms are independently drawn from a fixed distribution. While this shares the challenge of choosing from infinitely many arms, their setting assumes no regularity across arms. Therefore, their techniques and guarantees are not applicable to our setting.
@@ 1.3 Notations
For any integer $n$, we write $[\![n]\!]=\{1,\dots,n\}$. For a filtration $(\mathcal{F}_{t})_{t}$, we denote the conditional expectation by $\mathbb{E}_{t}\left[\cdot\right]=\mathbb{E}\left[\cdot\,|\,\mathcal{F}_{t}\right]$ (the underlying filtration will be clear from context). For real numbers $a$ and $b$, we write $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
@@ 2 Problem setting
Without loss of generality, we assume that the arm space is the unit interval $[0,1]$. [^2] Extensions to arbitrary metric spaces are discussed in Appendix G. An oblivious adversary selects a sequence of $T$ mean reward functions $(\mu_{t})_{t=1}^{T}$, where each $\mu_{t}:[0,1]\to[0,1]$ satisfies the following Lipschitz condition:
@@ 2.1 Significant regret and significant shifts
An arm $x\in[0,1]$ incurs significant regret on interval $[s_{1},s_{2}]$ if its cumulative regret on this interval is lower bounded as
@@
We call such arm an unsafe arm on interval $[s_{1},s_{2}]$, and otherwise we call it safe on this interval.
The right hand side of Definition 1 corresponds to the minimax-optimal regret for a stationary Lipschitz bandit over horizon $s_{2}-s_{1}$, and thus captures the level of regret that is statistically detectable. Conversely, if no such interval exists where arm $x$ incurs significant regret, then its cumulative regret over the entire horizon must remain small. A significant shift occurs when all arms become unsafe, i.e. when each arm has incurred significant regret over some interval.
@@
Let $\tau_{0}=1$. For $i\geq 1$, we define the $i^{\text{th}}$ significant shift as the smallest round $\tau_{i}\in]\tau_{i-1},T]$ such that for all arm $x\in[0,1]$, there exists an interval $[s_{1},s_{2}]\subseteq[\tau_{i-1},\tau_{i}]$ on which arm $x$ incurs significant regret (Definition 1). For all $i$, we call $[\tau_{i},\tau_{i+1}[$ a significant phase, and we denote by $\tilde{L}_{T}$ the (unknown) number of significant phases.
@@ 3 Algorithm MDBE: Multi-Depth Bin Elimination
Our algorithm operates in stable episodes, each split into blocks of doubling length with varying discretization levels. Within each block, a variant of the Successive Elimination algorithm removes empirically suboptimal regions. Since discarded bins may become optimal later, we use the replay mechanism of Suk and Kpotufe (2022) to revisit them and avoid missing promising regions. A key challenge is selecting the right discretization level for each replay. Our sampling procedure guarantees sufficient exploration at different scales, while a carefully designed bin eviction mechanism progressively eliminates regions that appear suboptimal based on empirical evidence.
@@
Episodes and blocks. The algorithm proceeds in episodes, indexed by $l\in\mathbb{N}$, with episode $l$ starting at round $t_{l}$ (with $t_{0}=1$). Each episode is divided into blocks, where the $m^{\text{th}}$ block lasts for $8^{m}$ rounds. Within the $m^{\text{th}}$ block, the action space $[0,1]$ is discretized into $2^{m}=(8^{m})^{1/3}$ bins, corresponding to the $m^{\text{th}}$ depth $\mathcal{T}_{m}$. We denote the block interval by $[\tau_{l,m},\tau_{l,m+1}[$, and maintain a set $\mathcal{B}_{\mathrm{MASTER}}(m)\subseteq\mathcal{T}_{m}$ of safe bins at depth $m$ (referred to as the MASTER set). A significant shift is detected when $\mathcal{B}_{\mathrm{MASTER}}(m)=\emptyset$, meaning all bins at depth $m$ have been evicted. In such cases, the block (and thus the episode) ends prematurely. Otherwise, a block ends naturally after $8^{m}$ rounds.
@@
      xChoose $x_{t}$ using Sampling scheme (Algorithm 2);
@@
Replays. To detect non-stationarity at different scales, the algorithm performs replays at various depths. Each block $[\tau_{l,m},\tau_{l,m+1}[$ always initiates a replay at depth $m$ by setting $\mathcal{B}_{\mathrm{MASTER}}(m)\leftarrow\mathcal{T}_{m}$. Additionally, replays may be triggered for all rounds $s=\tau_{l,m}+1,\dots,\tau_{l,m}+8^{m}-1$ at coarser depths $d<m$ with probability
@@
Each replay at depth $d$ runs for $8^{d}$ rounds, unless it is interrupted by the end of the current block. At the beginning of such a replay, the set of active bins at each active depth $d\in\mathcal{D}_{t}$ is reset to contain all bins at that depth: $\mathcal{B}_{t}(d)\leftarrow\mathcal{T}_{d}$ for all $d\in\mathcal{D}_{t}$. This particularly implies that active bins at the finest depth $m$ are also reset, i.e. $\mathcal{B}_{t}(m)\leftarrow\mathcal{T}_{m}$. However, the set $\mathcal{B}_{\mathrm{MASTER}}(m)$ is not reset and continues to reflect the outcome of eliminations from previous rounds within the block. When the replay at depth $d$ concludes, depth $d$ is simply removed from the set of active depths $\mathcal{D}_{t}$, but any other active depths may continue their replays independently. Thanks to this design, multiple replays at different depths can overlap in time, allowing the algorithm to monitor for potential significant shifts at multiple scales simultaneously (Algorithm 1).
Sampling scheme. At each round, the algorithm selects an action using a hierarchical, top-down sampling procedure from the minimum active depth $d_{0}(t)=\min\mathcal{D}_{t}$ down to the deepest active depth that contains an active bin (Algorithm 2). Remark that if $\mathcal{D}_{t}=\{m\}$, then $\mathcal{B}_{t}(m)=\mathcal{B}_{\mathrm{MASTER}}(m)$. In this case, the sampling procedure reduces to selecting a bin uniformly from $\mathcal{B}_{\mathrm{MASTER}}(m)$, and subsequently sampling an action from the chosen bin.
@@
Bin eviction criterion. Our algorithm aims to identify and eliminate arms that incur significant regret, as defined in Definition 1. However, since the algorithm only relies on bin estimates $\delta_{t}(B)$, we introduce a notion of significant regret for bins to make localized decisions.
@@
We say a bin $B\in\mathcal{T}_{d}$ incurs significant regret on interval $[s_{1},s_{2}]$ if
@@
The right-hand side of the inequality corresponds to the rate of the minimax regret of a $2^{d}$-armed bandit over interval $[s_{1},s_{2}]$. Intuitively, this threshold captures whether the cumulative sub-optimality of the bin is statistically significant over $[s_{1},s_{2}]$. Importantly, we will later show that if a bin meets this threshold, then all arms within the bin also suffer significant regret in the sense of Definition 1 (see Proposition 1). We emphasize that the notion of significant shifts (Definition 2) is defined independently of any depth (and thus independently of the depth used by the algorithm).
@@
where $\mathcal{F}_{t}$ is the natural filtration generated by the algorithm. Under uniform sampling within bin $B$, $\hat{\mu}_{t}(B)$ is an unbiased estimate of $\mu_{t}(B)$. In particular, during a replay at depth $d$, the algorithm leverages these estimates to potentially evict bins across active depths $d^{\prime}\geq d$ that appear suboptimal: over some $[s_{1},s_{2}]$ of length at most $8^{d}$ ($d$ is active), bin $B$ is evicted if
@@
where $c_{0}$ is a positive universal constant [^4] An exact value of $c_{0}$ can be derived from the analysis., and $\mathcal{B}_{[s_{1},s_{2}]}(d)$ denotes the set of bins that remain active throughout interval $[s_{1},s_{2}]$. Importantly, when a bin $B\in\mathcal{T}_{d}$ is evicted based on this criterion, it is immediately removed from the current set of active bins $\mathcal{B}_{t}$, and also permanently removed from the global MASTER set $\mathcal{B}_{\mathrm{MASTER}}(m)$. In addition, all children of $B$ in the dyadic tree are evicted as well. As a result, no future arms will be sampled from the entire region corresponding to $B$ and its descendants in $\mathcal{T}$ until a new replay is scheduled. The next proposition establishes that the eviction rule ($\star$3) is sound: evicted bins indeed correspond to areas where all arms suffer significant regret.
@@
If a bin $B\in\mathcal{T}_{d}$ incurs significant regret on an interval $[s_{1},s_{2}]$ with $s_{2}-s_{1}\leq 8^{d}$, then every point $x\in B$ also incurs significant regret over $[s_{1},s_{2}]$.

### v2

@@ 1 Introduction
Many practical applications involve decision-making over a continuous space of actions (or arms) where the environment evolves over time. For example, in dynamic pricing, the space of possible prices is inherently continuous, and market dynamics may shift abruptly. In such a setting, it is both natural and crucial to leverage smoothness in the reward function to generalize across similar actions, while also remaining adaptive to temporal changes. However, most existing works on non-stationary bandits focus on discrete action spaces and overlook the structure present in continuous domains.
@@
The non-stationary bandit problem extends the classical stochastic multi-armed bandit setting [Slivkins et al., 2019, Lattimore and Szepesvári, 2020] by allowing the mean reward functions to change over time. Over the past two decades, several frameworks have been proposed to quantify non-stationarity. Common approaches quantify the number of changes $L_{T}$ of any arm’s mean reward over $T$ rounds [Auer et al., 2002b, Garivier and Moulines, 2011], or their total variation $V_{T}$ in the reward functions [Besbes et al., 2014, Cheung et al., 2019]. Algorithms that are built upon these approaches achieve dynamic regret bounds of order $\sqrt{L_{T}T}$ or $V_{T}^{1/3}T^{2/3}$ respectively. Importantly, they all assume prior knowledge of the degree of non-stationarity $L_{T}$ or $V_{T}$. It remained an open question whether such rates could be achieved without knowing $L_{T}$ or $V_{T}$, until a line of work answered this affirmatively [Auer et al., 2019a, b, Chen et al., 2019, Wei and Luo, 2021].
@@
Recent works have introduced more refined measures of non-stationarity. In fact, the standard metrics $L_{T}$ and $V_{T}$ can be overly pessimistic: both of them may be arbitrarily large even when the best arm remains unchanged. To address this, Abbasi-Yadkori et al. [2023] designed an algorithm that tracks the number of times the best arm changes over $T$ rounds, $S_{T}$, and showed that it achieves $\sqrt{S_{T}T}$ for the dynamic regret, adaptively [^1] In all our work, ”adaptively” should be understood the same as ”without knowledge of the non-stationarity”.. The important work of Suk and Kpotufe [2022] proposed a finer notion of non-stationarity called significant shifts, which identifies only changes that meaningfully affect regret through large aggregate differences in reward. In the $K$-armed bandit setting, their algorithm achieves the optimal rate of $\sqrt{\tilde{L}_{T}T}$ where $\tilde{L}_{T}$ is the number of significant shifts over $T$ rounds, with $\tilde{L}_{T}\ll L_{T}$.
@@
In this work, we extend the ideas of Suk and Kpotufe [2022] to the setting of Lipschitz bandits, where the arm space is continuous (typically the interval $[0,1]$) and the reward functions satisfy a Lipschitz condition at each round. Despite the maturity of research in both non-stationary and Lipschitz bandits individually, their intersection remains unexplored. Our work is the first one to tackle the problem of non-stationary Lipschitz bandits.
@@ 1.1 Outline and contributions
Discretization level. In the stationary Lipschitz bandit setting, a common approach is to discretize the continuous arm space into a finite set of bins, thereby allowing the application of standard multi-armed bandit algorithms [Kleinberg, 2004]. When the time horizon $T$ is known, the discretization level can be carefully calibrated, typically using $\mathcal{O}(T^{1/3})$ bins. This choice reflects the optimal balance between discretization error and the regret incurred by running a bandit algorithm over the discretized arms, leading to the minimax-optimal regret of $\mathcal{O}(T^{2/3})$. In the non-stationary setting, however, the durations of the time intervals during which the mean reward remains stable are unknown and not rigorously defined. Thus, direct horizon-dependent tuning becomes infeasible. This raises a key question: how to adaptively adjust the discretization level to match the unknown degree of non-stationarity in the environment?
Forced exploration via restarts and replays. In non-stationary bandits, detecting changes in the reward distribution requires frequent exploration of suboptimal arms. In the $K$-armed bandit setting, this is often handled through periodic restarts, where all $K$ arms are re-explored to detect shifts of the reward function. In the continuous setting, this strategy is significantly harder: not only must we decide when to restart, but also at what resolution to discretize the arm space during such restarts. Finer discretization enables more precise detection of changes but significantly increases the cost of exploration. Balancing this trade-off is a central technical challenge tackled in our work.
This paper is organized as follows. In Section˜2, we formally introduce the problem setting, and present our algorithm, MDBE, in Section˜3. Our main results are stated in Section˜4, followed by a high-level proof sketch of its regret analysis in Section˜5. For completeness, the full pseudocode of MDBE is provided in Appendix˜B, and all detailed regret proofs can be found in Appendix˜D.
@@ 1.2 Related Work
Non-stationary bandits beyond $K$-armed. A growing body of work has extended non-stationary bandit models beyond the classical $K$-armed setting. These include linear and generalized linear bandits [Russac et al., 2019, Faury et al., 2021], contextual bandits [Suk and Kpotufe, 2021], convex bandits [Zhao et al., 2021, Wang, 2023, Liu et al., 2025], kernelized bandits [Hong et al., 2023, Iwazaki and Takeno, 2025, Cai and Scarlett, 2025], and more structured problems [Seznec et al., 2020, Vernade et al., 2020, Azizi et al., 2022]. Some works further explore temporal regularity assumptions, such as smooth or slowly varying reward functions [Slivkins and Upfal, 2008, Krishnamurthy and Gopalan, 2021, Jia et al., 2023]. Across these settings, non-stationarity is typically quantified using either the number of global changes $L_{T}$ or the total variation $V_{T}$.
@@
Bandits with infinite arms. In Lipschitz bandits [Agrawal, 1995], the infinite arm space is tackled via discretization strategies, either fixed or adaptive, that exploit the Lipschitz continuity of the reward function [Kleinberg, 2004, Kleinberg et al., 2008, Bubeck et al., 2011b, a, Magureanu et al., 2014]. More recently, adversarial formulations have been studied, where the goal is to minimize the regret in hindsight [Podimata and Slivkins, 2021, Kang et al., 2023b]. These differ fundamentally from our focus on dynamic regret in a stochastic setting. A separate line of work investigates infinite-armed bandits in the so-called reservoir model, where arm means are drawn independently from a fixed distribution [Wang et al., 2008, Carpentier and Valko, 2015, De Heide et al., 2021, Kim et al., 2022, 2024]. Unlike Lipschitz bandits, these models assume no structural relationship between arms and therefore do not exploit similarity across them. Hence, direct comparisons to our setting do not apply. The only work we are aware of that addresses the problem of non-stationary Lipschitz bandits is Kang et al. [2023a], who propose a zooming-like algorithm to tune a generalized bandit framework. However, their formulation of non-stationary Lipschitz bandits is tailored to their specific setting and does not achieve optimal theoretical guarantees.
@@
Tracking significant shifts. The idea of tracking significant shifts was first introduced in the $K$-armed setting by Suk and Kpotufe [2022], and has since been extended to contextual bandits [Suk and Kpotufe, 2023], preference-based settings [Suk and Agarwal, 2023, Buening and Saha, 2023, Suk and Agarwal, 2024], and smooth non-stationary models with temporal structure [Suk, 2024].
@@
The works closest to ours are Suk and Kpotufe [2023] and Suk and Kim [2025], both extending the framework of tracking significant shifts beyond the classical $K$-armed setting. The former considers contextual bandits where rewards are Lipschitz in the context space, but with a finite number of arms. After discretizing the context space, their algorithm effectively reduces to a finite-arm bandit problem at each context, allowing the use of techniques from the $K$-armed setting [Suk and Kpotufe, 2022] without requiring discretization of the arm space. The latter studies infinite-armed bandits under a reservoir model, where arms are independently drawn from a fixed distribution. While this shares the challenge of choosing from infinitely many arms, their setting assumes no regularity across arms. Therefore, their techniques and guarantees are not applicable to our setting.
@@ 1.3 Notations
For any integer $n$, we write $n]\!]=\{1,\dots,n\}$. For a filtration $(\mathcal{F}_{t})_{t}$, we denote the conditional expectation by $\mathbb{E}_{t}\left[\cdot\right]=\mathbb{E}\left[\cdot\,|\,\mathcal{F}_{t}\right]$ (the underlying filtration will be clear from context). For real numbers $a$ and $b$, we write $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
@@ 2 Problem setting
Without loss of generality, we assume that the arm space is the unit interval [0,1][0,1]. [^2] Extensions to arbitrary metric spaces and other Lipschitz constants are discussed in Appendix G. An oblivious adversary selects a sequence of $T$ mean reward functions $(\mu_{t})_{t=1}^{T}$, where each $\mu_{t}:[0,1]\to[0,1]$ satisfies the following Lipschitz condition:
@@ 2.1 Significant regret and significant shifts
An arm $x\in\mathcal{X}$ incurs significant regret on interval $s_{1},s_{2}]$ if its cumulative regret on this interval is lower bounded as
@@
We call such arm an unsafe arm on interval $s_{1},s_{2}]$, and otherwise we call it safe on this interval.
The right hand side of Definition˜1 corresponds to the minimax-optimal regret for a stationary Lipschitz bandit over horizon $s_{2}-s_{1}$, and thus captures the level of regret that is statistically detectable. Conversely, if no such interval exists where arm $x$ incurs significant regret, then its cumulative regret over the entire horizon must remain small. A significant shift occurs when all arms become unsafe, i.e. when each arm has incurred significant regret over some interval.
@@
Let $\tau_{0}=1$. For $i\geq 1$, we define the $i^{\text{th}}$ significant shift as the smallest round $\tau_{i}\in]\tau_{i-1},T]$ such that for all arm $x\in\mathcal{X}$, there exists an interval $s_{1},s_{2}]\subseteq[\tau_{i-1},\tau_{i}]$ on which arm $x$ incurs significant regret (Definition˜1). For all $i$, we call $\tau_{i},\tau_{i+1}[$ a significant phase, and we denote by $\tilde{L}_{T}$ the (unknown) number of significant phases.
@@
For example, consider a recommendation system with a continuous pool of content (e.g. movies), indexed by $x\in[0,1]$, where nearby values of $x$ correspond to similar content types. Suppose there exist two regions of high and comparable user preference, centred around $x_{1}=0.3$ and $x_{2}=0.7$. Imagine a scenario where preferences near $x_{1}$ remain stable over time (e.g. a timeless classic), while preferences near $x_{2}$ undergo very frequent changes (e.g. a trending topic that evolves daily). In this case, an algorithm that consistently recommends content near $x_{1}$ would incur little to no regret, even though the underlying reward function changes frequently in other regions. From a global perspective, the number of changes or the total variation in mean reward could be as large as $L_{T}=V_{T}=\mathcal{O}(T)$. An algorithm that relies solely on such metrics would unnecessarily restart its estimates too frequently, as it would overestimate the effective difficulty of the problem. This leads to both theoretical sub-optimality and practical inefficiency. However, such changes do not constitute a significant shift, as the latter captures only changes that are statistically consequential to learning, rather than indiscriminately counting all shifts in the environment.
@@ 3 Algorithm MDBE: Multi-Depth Bin Elimination
Our algorithm operates in stable episodes, each split into blocks of doubling length with varying discretization levels. Within each block, a variant of the Successive Elimination algorithm removes empirically suboptimal regions. Since discarded bins may become optimal later, we use the replay mechanism of Suk and Kpotufe [2022] to revisit them and avoid missing promising regions. A key challenge is selecting the right discretization level for each replay. Our sampling procedure guarantees sufficient exploration at different scales, while a carefully designed bin eviction mechanism progressively eliminates regions that appear suboptimal based on empirical evidence.
@@
Episodes and blocks. The algorithm proceeds in episodes, indexed by $l\in\mathbb{N}$, with episode $l$ starting at round $t_{l}$ (with $t_{0}=1$). Each episode is divided into blocks, where the $m^{\text{th}}$ block lasts for $8^{m}$ rounds. Within the $m^{\text{th}}$ block, the action space $[0,1]$ is discretized into $2^{m}=(8^{m})^{1/3}$ bins, corresponding to the $m^{\text{th}}$ depth $\mathcal{T}_{m}$. We denote the block interval by $\tau_{l,m},\tau_{l,m+1}[$, and maintain a set $\mathcal{B}_{\mathrm{MASTER}}(m)\subseteq\mathcal{T}_{m}$ of safe bins at depth $m$ (referred to as the MASTER set). A significant shift is detected when $\mathcal{B}_{\mathrm{MASTER}}(m)=\emptyset$, meaning all bins at depth $m$ have been evicted. In such cases, the block (and thus the episode) ends prematurely. Otherwise, a block ends naturally after $8^{m}$ rounds.
@@
 xChoose $x_{t}$ using Sampling scheme (Algorithm˜2);
@@
Replays. To detect non-stationarity at different scales, the algorithm performs replays at various depths. Each block $\tau_{l,m},\tau_{l,m+1}[$ always initiates a replay at depth $m$ by setting $\mathcal{B}_{\mathrm{MASTER}}(m)\leftarrow\mathcal{T}_{m}$. Additionally, replays may be triggered for all rounds $s=\tau_{l,m}+1,\dots,\tau_{l,m}+8^{m}-1$ at coarser depths $d<m$ with probability
@@
Each replay at depth $d$ runs for $8^{d}$ rounds, unless it is interrupted by the end of the current block. At the beginning of such a replay, the set of active bins at each active depth $d\in\mathcal{D}_{t}$ is reset to contain all bins at that depth: $\mathcal{B}_{t}(d)\leftarrow\mathcal{T}_{d}$ for all $d\in\mathcal{D}_{t}$. When the replay at depth $d$ concludes, depth $d$ is simply removed from the set of active depths $\mathcal{D}_{t}$, but any other active depths may continue their replays independently. Thanks to this design, multiple replays at different depths can overlap in time, allowing the algorithm to monitor for potential significant shifts at multiple scales simultaneously (Algorithm˜1).
Sampling scheme. At each round, the algorithm selects an action using a hierarchical, top-down sampling procedure from the minimum active depth $d_{0}(t)=\min\mathcal{D}_{t}$ down to the deepest active depth that contains an active bin (Algorithm˜2). Remark that if $\mathcal{D}_{t}=\{m\}$, then $\mathcal{B}_{t}(m)=\mathcal{B}_{\mathrm{MASTER}}(m)$. In this case, the sampling procedure reduces to selecting a bin uniformly from $\mathcal{B}_{\mathrm{MASTER}}(m)$, and subsequently sampling an action from the chosen bin (see Figure˜1).
@@
Bin eviction criterion. Our algorithm aims to identify and eliminate arms that incur significant regret, as defined in Definition˜1. However, since the algorithm only relies on bin estimates $\delta_{t}(B)$, we introduce a notion of significant regret for bins to make localized decisions.
@@
We say a bin $B\in\mathcal{T}_{d}$ incurs significant regret on interval $s_{1},s_{2}]$ if
@@
The right-hand side of the inequality corresponds to the rate of the minimax regret of a $2^{d}$-armed bandit over interval $s_{1},s_{2}]$. Intuitively, this threshold captures whether the cumulative sub-optimality of the bin is statistically significant over $s_{1},s_{2}]$. Importantly, we will later show that if a bin meets this threshold, then all arms within the bin also suffer significant regret in the sense of Definition˜1 (see Proposition˜1). We emphasize that the notion of significant shifts (Definition˜2) is defined independently of any depth (and thus independently of the depth used by the algorithm).
@@
where $\mathcal{F}_{t}$ is the natural filtration generated by the algorithm. Under uniform sampling within bin $B$, $\hat{\mu}_{t}(B)$ is an unbiased estimate of $\mu_{t}(B)$. In particular, during a replay at depth $d$, the algorithm leverages these estimates to potentially evict bins across active depths $d^{\prime}\geq d$ that appear suboptimal: over some $s_{1},s_{2}]$ of length at most $8^{d}$ ($d$ is active), bin $B$ is evicted if
@@
where $c_{0}$ is a positive universal constant [^4] An exact value of $c_{0}$ can be derived from the analysis., and $\mathcal{B}_{[s_{1},s_{2}]}(d)$ denotes the set of bins that remain active throughout interval $s_{1},s_{2}]$. Importantly, when a bin $B\in\mathcal{T}_{d}$ is evicted based on this criterion, it is immediately removed from the current set of active bins $\mathcal{B}_{t}$, and also permanently removed from the global MASTER set $\mathcal{B}_{\mathrm{MASTER}}(m)$. In addition, all children of $B$ in the dyadic tree are evicted as well. As a result, no future arms will be sampled from the entire region corresponding to $B$ and its descendants in $\mathcal{T}$ until a new replay is scheduled. The next proposition establishes that the eviction rule ($\star$3) is sound: evicted bins indeed correspond to areas where all arms suffer significant regret.
@@
If a bin $B\in\mathcal{T}_{d}$ incurs significant regret on an interval $s_{1},s_{2}]$ with $s_{2}-s_{1}\leq 8^{d}$, then every point $x\in B$ also incurs significant regret over $s_{1},s_{2}]$.
Figure 1: Example of sampling with $m=4$; active bins are in blue. Left: At time $t_{1}$, depths 33 and $m$ are active. A sample path may select bin $B_{3,3}$ uniformly at random (u.a.r.) at depth $3$, then $B_{m,6}$ u.a.r. among its active children, then arm $x_{t}$ u.a.r. in $B_{m,6}$ (red path). Center: At $t_{1}+1$, a replay starts at depth $1$. A path may go through $B_{1,1}\rightarrow B_{3,2}\rightarrow B_{m,4}$, selecting $x_{t}$ in $B_{m,4}$ (red path). Alternatively, $B_{1,2}$ could be chosen; with no active children, $x_{t}$ is sampled directly from it (green choice). Right: At $t_{1}+9$, depth 11 exits replay. Bin has been $B_{3,2}$ eliminated during the replay, and a path may select $B_{3,1}\rightarrow B_{m,1}$, then $x_{t}$ in $B_{m,1}$ (red path).